OpenStudy (anonymous):

complete the ordered pair for the equation y=2x+5. ( ,-1)

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

no wait thats wrong

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

y=2x+5 substitute in -1 for y -1=2x+5 -5 -5 -6=2x divide by 2 -3=x (-3, -1)